# GradMix: Gradient-based Selective Mixup for Robust Data Augmentation in Class-Incremental Learning

### v1

@@ Abstract
In the context of continual learning, acquiring new knowledge while maintaining previous knowledge presents a significant challenge. Existing methods often use experience replay techniques that store a small portion of previous task data for training. In experience replay approaches, data augmentation has emerged as a promising strategy to further improve the model performance by mixing limited previous task data with sufficient current task data. However, we theoretically and empirically analyze that training with mixed samples from random sample pairs may harm the knowledge of previous tasks and cause greater catastrophic forgetting. We then propose GradMix, a robust data augmentation method specifically designed for mitigating catastrophic forgetting in class-incremental learning. GradMix performs gradient-based selective mixup using a class-based criterion that mixes only samples from helpful class pairs and not from detrimental class pairs for reducing catastrophic forgetting. Our experiments on various real datasets show that GradMix outperforms data augmentation baselines in accuracy by minimizing the forgetting of previous knowledge.
1 Introduction
@@
Humans can learn incrementally and accumulate knowledge as they encounter new experiences throughout their lifetime. In contrast, traditional machine learning models do not have the ability to update themselves after learning from the given data once. Continual learning, also known as lifelong learning DBLP:series/synthesis/2016Chen , refers to a paradigm in artificial intelligence where a model continuously learns new tasks over time while retaining the knowledge of previously-acquired tasks. Continual learning is becoming crucial for real-world applications where the environment or tasks change over time, such as robotics DBLP:journals/inffus/LesortLSMFR20 , autonomous driving DBLP:journals/nn/VerwimpYPHMPLT23 , personalized recommendations DBLP:conf/sc/XieRLYXWLAXS20 ; DBLP:conf/sigir/CaiZDD0TZ22 , and more.
@@
There are three main scenarios of continual learning DBLP:journals/corr/abs-1904-07734 : task-incremental, domain-incremental, and class-incremental learning. In this paper, we focus on class-incremental learning, which is the most challenging scenario where a model incrementally learns new classes of data over tasks. A key challenge for class-incremental learning is catastrophic forgetting MCCLOSKEY1989109 , where a model forgets knowledge of previously-learned classes while learning new classes. Hence, balancing the trade-off between retaining previous knowledge and adapting to new information is important to achieve effective class-incremental learning ABRAHAM200573 ; 10.3389/fpsyg.2013.00504 ; DBLP:conf/cvpr/KimH23 .
@@
Several approaches have been proposed to mitigate catastrophic forgetting in class-incremental learning. One simple and effective approach is experience replay DBLP:journals/corr/abs-1902-10486 ; DBLP:conf/nips/RolnickASLW19 , which stores a subset of data from previous tasks in a buffer and uses it for training with the recently-arrived current task data. As the performance of experience replay is better than other buffer-free approaches DBLP:conf/eccv/LiH16 ; DBLP:journals/corr/KirkpatrickPRVD16 ; DBLP:conf/icml/ZenkePG17 , it is now a standard basis method and widely used with other approaches in the framework of class-incremental learning DBLP:conf/cvpr/RebuffiKSL17 ; DBLP:conf/nips/BuzzegaBPAC20 ; DBLP:conf/cvpr/YanX021 .
@@
However, experience replay does not fully solve the problem of catastrophic forgetting. Since there is a constraint on the buffer size in the continual learning literature due to memory capacity, computational efficiency, and privacy issues DBLP:journals/pami/LangeAMPJLST22 ; DBLP:journals/pami/WangZSZ24 , we can only store a small portion of data for previous tasks in a buffer. This memory constraint leads to an imbalance between the previous task data and the current task data, where there is a large amount of current task data compared to a small amount of buffer data for previous tasks. After training with the imbalanced data at each task, the model performance of previous tasks is much lower than that of the current task.
@@
One solution to address the problem of limited buffer data is data augmentation. For image datasets, image transformations like random rotation, flipping, and cropping can be used to increase the proportion of buffer data. However, transforming images still has limitations in improving the diversity of buffer data because the source buffer size is already too small. Here mixup-based data augmentation methods may be better to improve the diversity of buffer data by mixing them with a large amount of current task data. In practice, several works empirically show that applying mixup-based methods to class-incremental learning can further improve the model performance DBLP:conf/cvpr/MiKLYF20 ; DBLP:conf/cvpr/BangKY0C21 . However, these works overlook catastrophic forgetting while applying mixup-based methods, focusing only on their augmentation and generalization capabilities. We theoretically and empirically find that applying vanilla Mixup can worsen catastrophic forgetting, where the cause is an interference between the average gradient vectors of augmented training data and previous task data when they are in the opposite direction.
@@
Figure 1: Overview of GradMix. The gradients of the original samples and the mixed samples are denoted by $g$ and $\tilde{g}$, respectively. Given the average gradient of buffer data (denoted as $\overline{g}_{buf}$), GradMix performs gradient-based selective mixup to mitigate catastrophic forgetting by mixing only samples from helpful class pairs (e.g., ‘Dog’–‘Deer’, where $\theta\leq 90^{\circ}$) and not from detrimental class pairs (e.g., ‘Dog’–‘Automobile’, where $\theta>90^{\circ}$).
@@ 2 Related Work
Class-incremental learning is a challenging scenario within the field of continual learning, focusing on incrementally learning new classes of data over time. The main problem in class-incremental learning is catastrophic forgetting MCCLOSKEY1989109 , where a model forgets knowledge about previously-learned classes while acquiring new classes. Existing continual learning techniques can be categorized as experience replay, parameter regularization, and dynamic networks. Experience replay methods store and replay a subset of previous data to explicitly mitigate forgetting DBLP:journals/corr/abs-1902-10486 ; DBLP:conf/nips/AljundiBTCCLP19 ; DBLP:conf/aaai/ShimMJSKJ21 ; DBLP:conf/cvpr/MaiLKS21 . Parameter regularization methods add constraints on model parameters to prevent drastic changes that could lead to forgetting DBLP:journals/corr/KirkpatrickPRVD16 ; DBLP:conf/icml/ZenkePG17 ; DBLP:conf/eccv/AljundiBERT18 ; DBLP:conf/iclr/KolouriKSP20 . Dynamic network methods dynamically expand or adapt the network architecture to learn new tasks without overwriting prior knowledge DBLP:conf/cvpr/YanX021 ; DBLP:conf/eccv/WangZYZ22 ; DBLP:conf/iclr/0001WYZ23 . There are also other lines of research including knowledge distillation DBLP:conf/eccv/LiH16 ; DBLP:conf/cvpr/RebuffiKSL17 ; DBLP:conf/cvpr/HouPLWL19 and model rectification DBLP:conf/cvpr/WuCWYLGF19 ; DBLP:conf/cvpr/ZhaoXGZX20 . Among these categories, experience replay is widely used in the class-incremental learning literature, and it is often combined with other techniques from different categories DBLP:conf/nips/BuzzegaBPAC20 ; DBLP:conf/cvpr/YanX021 . In this paper, we focus on the limitation of experience replay methods where the amount of stored buffer data is limited and propose a robust data augmentation method that minimizes the forgetting of previous tasks.
@@
Data augmentation is an effective method to enhance the generalization and robustness of models by generating additional training data. Traditional data augmentation techniques include image transformations such as random rotation, flipping, and cropping, which introduce variations to the input images DBLP:journals/jbd/ShortenK19 ; DBLP:journals/corr/abs-2204-08610 . Recent works propose an automatic strategy to find an effective data augmentation policy DBLP:conf/cvpr/CubukZMVL19 ; DBLP:conf/nips/CubukZS020 ; DBLP:conf/iccv/MullerH21 . However, these policy-based approaches do not consider continual learning scenarios, which may result in severe catastrophic forgetting as we show in our experiments.
@@
Another line of research is Mixup DBLP:conf/iclr/ZhangCDL18 , which generates new synthetic data by linearly interpolating between pairs of features and their corresponding labels, and there are variant mixup-based methods DBLP:conf/icml/VermaLBNMLB19 ; DBLP:conf/iccv/YunHCOYC19 ; harris2021fmixenhancingmixedsample . In addition, imbalance-aware mixup methods have been proposed to improve model robustness on imbalanced datasets DBLP:conf/eccv/ChouCPWJ20 ; DBLP:conf/miccai/GaldranCB21 . Furthermore, selective mixup is introduced by selecting sample pairs for mixup based on specific criteria instead of randomly, and it improves the model’s ability for domain adaptation DBLP:conf/aaai/XuZNLWTZ20 and out-of-distribution (OOD) generalization DBLP:conf/icml/Yao0LZL0F22 .
@@
A recent line of research applies existing data augmentation methods in class-incremental learning to further improve accuracy. Most existing works simply combine experience replay methods with vanilla Mixup DBLP:conf/cvpr/MiKLYF20 ; DBLP:conf/nips/KumariW0B22 , RandAugment DBLP:conf/nips/ZhangPFBLJ22 , or a combination of existing data augmentation methods DBLP:conf/cvpr/BangKY0C21 . Extending this line of work, DBLP:conf/aaai/QiangH0L0Z23 improves the sampling and mixing factor strategies in Mixup to address distribution discrepancies between the imbalanced training set and the balanced test set in continual learning, which is similar to existing imbalance-aware mixup methods DBLP:conf/eccv/ChouCPWJ20 ; DBLP:conf/miccai/GaldranCB21 . Another work DBLP:conf/nips/KimXKK022 addresses class-incremental learning by using out-of-distribution detection methods that utilize image rotation as a data augmentation technique. Another line of research is using generative models to generate data for previous tasks DBLP:conf/nips/ShinLKK17 ; DBLP:conf/icml/GaoL23a ; DBLP:conf/cvpr/KimCKTB24 , but this approach incurs additional costs for continuously training generative models across tasks. In addition, there is a work DBLP:conf/nips/ZhuCZL21 to achieve comparable performance with experience replay methods without storing buffer data by utilizing class augmentation and semantic augmentation techniques. However, all these works simply use existing data augmentation techniques for further improvements and do not propose any novel data augmentation methods specifically designed for continual learning scenarios. In comparison, GradMix is a robust data augmentation method designed to minimize catastrophic forgetting in class-incremental learning. We leave more detailed comparisons with these works in Sec. D.
3 Preliminaries
@@
We set notations for class-incremental learning. We denote each task as $T_{l}=\{d_{i}=(x_{i},y_{i})\}_{i=1}^{k}$, where $x$, $y$, and $k$ represent the feature, the true label, and the number of samples, respectively. We assume disjoint sets of classes as tasks, where there are no overlapping classes between different tasks. The sequence of tasks is given by $T=\{T_{1},\ldots,T_{L}\}$, where $L$ is the total number of tasks. At the $l^{th}$ step, we consider $T_{l}$ as the current task and $\{T_{1},\ldots,T_{l-1}\}$ as previous tasks. Let the set of previous classes be $\mathbb{Y}_{p}$ and the set of current classes be $\mathbb{Y}_{c}$. We store a small subset of data for the previous tasks in a buffer to employ experience replay during training. We denote the subset of each previous task as $\mathcal{M}_{t}=\{d_{j}=(x_{j},y_{j})\}_{j=1}^{m}$, where $m$ is the number of samples. The buffer data for the previous tasks is then represented as =$\mathcal{M}=\{\mathcal{M}_{1},\ldots,\mathcal{M}_{l-1}\}$. Since we cannot store large buffer data due to the possible problems of memory, efficiency, and privacy during continual learning, we assume the buffer size is small (i.e., $m\ll k$) DBLP:journals/corr/abs-1902-10486 .
@@
Catastrophic forgetting is the main challenge of continual learning, where the performance on previously-learned tasks degrades while adapting to the current task. The condition for the occurrence of catastrophic forgetting is theoretically derived using the gradients of previous and current data with respect to the model DBLP:conf/nips/Lopez-PazR17 ; DBLP:conf/iclr/ChaudhryRRE19 ; DBLP:conf/nips/AljundiLGB19 . At the $l^{th}$ step, let $G_{y}$ be a group of samples whose label is $y\in\mathbb{Y}_{p}$, and $T_{l}$ be the current task data used for training. We can approximate the average loss of $G_{y}$ after model training with $T_{l}$ by employing first-order Taylor series approximation as follows:
~$\tilde{\ell}(f_{\theta},G_{y})=\ell(f_{\theta}^{l-1},G_{y})-\eta\nabla_{\theta}\ell(f_{\theta}^{l-1},G_{y})^{\top}\nabla_{\theta}\ell(f_{\theta}^{l-1},T_{l}),$ (1)
where $\tilde{\ell}(\cdot,\cdot)$ is the approximated average cross-entropy loss, $\ell(\cdot,\cdot)$ is the exact average cross-entropy loss, $f_{\theta}^{l-1}$ is a previous model, $f_{\theta}$ is an updated model, and $\eta$ is a learning rate. Hence, the inner product of the two average gradient vectors of the previous and current data determines whether catastrophic forgetting occurs or not. If the inner product value is negative, it leads to the increase of the loss for $G_{y}$ after training, which implies that catastrophic forgetting occurs for $G_{y}$ while learning $T_{l}$, and vice versa.
@@
Mixup DBLP:conf/iclr/ZhangCDL18 is a data augmentation method that generates new training samples by combining random pairs of existing samples of features and their labels. Given two arbitrary samples $d_{i}=(x_{i},y_{i})$ and $d_{j}=(x_{j},y_{j})$ from the dataset, a new mixed sample $\tilde{d}_{ij}=(\tilde{x}_{ij},\tilde{y}_{ij})$ is generated from the two original samples as follows:
$\tilde{x}_{ij}=\lambda x_{i}+(1-\lambda)x_{j},\ \tilde{y}_{ij}=\lambda y_{i}+(1-\lambda)y_{j},$ (2)
@@
4 Catastrophic Forgetting in Mixup
In this section, we demonstrate that simply applying Mixup to experience replay methods in class-incremental learning can lead to greater catastrophic forgetting of previously learned classes. In this paper, we use a simple yet representative experience replay method, ER DBLP:journals/corr/abs-1902-10486 , as a basis method for class-incremental learning, although other methods could also be used.
4.1 Theoretical Analysis

### v2

@@ Abstract.
In the context of continual learning, acquiring new knowledge while maintaining previous knowledge presents a significant challenge. Existing methods often use experience replay techniques that store a small portion of previous task data for training. In experience replay approaches, data augmentation has emerged as a promising strategy to further improve the model performance by mixing limited previous task data with sufficient current task data. However, we theoretically and empirically analyze that training with mixed samples from random sample pairs may harm the knowledge of previous tasks and cause greater catastrophic forgetting. We then propose GradMix, a robust data augmentation method specifically designed for mitigating catastrophic forgetting in class-incremental learning. GradMix performs gradient-based selective mixup using a class-based criterion that mixes only samples from helpful class pairs and not from detrimental class pairs for reducing catastrophic forgetting. Our experiments on various real datasets show that GradMix outperforms data augmentation baselines in accuracy by minimizing the forgetting of previous knowledge. The source code is available at: https://github.com/minsu716-kim/GradMix.
Continual learning, Class-incremental learning, Data augmentation
journalyear: 2026conference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreaisbn: 979-8-4007-2258-5/2026/08doi: 10.1145/XXXXXXX.XXXXXXXccs: Computing methodologies Machine learning
1. Introduction
@@
Humans can learn incrementally and accumulate knowledge as they encounter new experiences throughout their lifetime. In contrast, traditional machine learning models do not have the ability to update themselves after learning from the given data once. Continual learning, also known as lifelong learning (Chen and Liu, 2016), refers to a paradigm in artificial intelligence where a model continuously learns new tasks over time while retaining the knowledge of previously-acquired tasks. Continual learning is becoming crucial for real-world applications where the environment or tasks change over time, such as robotics (Lesort et al., 2020), autonomous driving (Verwimp et al., 2023), personalized recommendations (Xie et al., 2020; Cai et al., 2022), and more.
@@
There are three main scenarios of continual learning (van de Ven and Tolias, 2019): task-incremental, domain-incremental, and class-incremental learning. In this paper, we focus on class-incremental learning, which is the most challenging scenario where a model incrementally learns new classes of data over tasks. A key challenge for class-incremental learning is catastrophic forgetting (McCloskey and Cohen, 1989), where a model forgets knowledge of previously-learned classes while learning new classes. Hence, balancing the trade-off between retaining previous knowledge and adapting to new information is important to achieve effective class-incremental learning (Abraham and Robins, 2005; Mermillod et al., 2013; Kim and Han, 2023).
@@
Several approaches have been proposed to mitigate catastrophic forgetting in class-incremental learning. One simple and effective approach is experience replay (Chaudhry et al., 2019b; Rolnick et al., 2019), which stores a subset of data from previous tasks in a buffer and uses it for training with the recently-arrived current task data. As the performance of experience replay is better than other buffer-free approaches (Li and Hoiem, 2016; Kirkpatrick et al., 2016; Zenke et al., 2017), it is now a standard basis method and widely used with other approaches in the framework of class-incremental learning (Rebuffi et al., 2017; Buzzega et al., 2020; Yan et al., 2021).
@@
However, experience replay does not fully solve the problem of catastrophic forgetting. Since there is a constraint on the buffer size in the continual learning literature due to memory capacity, computational efficiency, and privacy issues (Lange et al., 2022; Wang et al., 2024), we can only store a small portion of data for previous tasks in a buffer. This memory constraint leads to an imbalance between the previous task data and the current task data, where there is a large amount of current task data compared to a small amount of buffer data for previous tasks. After training with the imbalanced data at each task, the model performance of previous tasks is much lower than that of the current task.
@@
One solution to address the problem of limited buffer data is data augmentation. For image datasets, transformations like random rotation, flipping, and cropping can be used to increase the proportion of buffer data. However, transforming images still has limitations in improving the diversity of buffer data because the source buffer size is already too small. Here mixup-based data augmentation methods may be better to improve the diversity of buffer data by mixing them with a large amount of current task data. In practice, several works empirically show that applying mixup-based methods to class-incremental learning can further improve the model performance (Mi et al., 2020; Bang et al., 2021). However, these works overlook catastrophic forgetting while applying mixup-based methods, focusing only on their augmentation and generalization capabilities. We theoretically and empirically find that applying vanilla Mixup can worsen catastrophic forgetting, where the cause is an interference between the average gradient vectors of augmented training data and previous task data when they are in the opposite direction.
@@
Figure 1. Overview of GradMix. The gradients of the original samples and the mixed samples are denoted by $g$ and $\tilde{g}$, respectively. Given the average gradient of buffer data (denoted as $\overline{g}_{buf}$), GradMix performs gradient-based selective mixup to mitigate catastrophic forgetting by mixing only samples from helpful class pairs (e.g., ‘Dog’–‘Deer’, where $\theta\leq 90^{\circ}$) and not from detrimental class pairs (e.g., ‘Dog’–‘Automobile’, where $\theta>90^{\circ}$).
@@ 2. Related Work
Class-incremental learning is a challenging scenario within the field of continual learning, focusing on incrementally learning new classes of data over time. The main problem in class-incremental learning is catastrophic forgetting (McCloskey and Cohen, 1989), where a model forgets knowledge about previously-learned classes while acquiring new classes. Existing continual learning techniques can be categorized as experience replay, parameter regularization, and dynamic networks. Experience replay methods store and replay a subset of previous data to explicitly mitigate forgetting (Chaudhry et al., 2019b; Aljundi et al., 2019a; Shim et al., 2021; Mai et al., 2021). Parameter regularization methods add constraints on model parameters to prevent drastic changes that could lead to forgetting (Kirkpatrick et al., 2016; Zenke et al., 2017; Aljundi et al., 2018; Kolouri et al., 2020). Dynamic network methods dynamically expand or adapt the network architecture to learn new tasks without overwriting prior knowledge (Yan et al., 2021; Wang et al., 2022; Zhou et al., 2023). There are also other lines of research including knowledge distillation (Li and Hoiem, 2016; Rebuffi et al., 2017; Hou et al., 2019) and model rectification (Wu et al., 2019; Zhao et al., 2020). Among these categories, experience replay is widely used in the class-incremental learning literature, and it is often combined with other techniques from different categories (Buzzega et al., 2020; Yan et al., 2021). In this paper, we focus on the limitation of experience replay methods where the amount of stored buffer data is limited and propose a robust data augmentation method that minimizes the forgetting of previous tasks.
@@
Data augmentation is an effective method to enhance the generalization and robustness of models by generating additional training data. Traditional data augmentation techniques include image transformations such as random rotation, flipping, and cropping, which introduce variations to the input images (Shorten and Khoshgoftaar, 2019; Yang et al., 2022). Recent works propose an automatic strategy to find an effective data augmentation policy (Cubuk et al., 2019, 2020; Müller and Hutter, 2021). However, these policy-based approaches do not consider continual learning scenarios, which may result in severe catastrophic forgetting as we show in experiments.
@@
Another line of research is Mixup (Zhang et al., 2018), which generates new synthetic data by linearly interpolating between pairs of features and their corresponding labels, and there are variant mixup-based methods (Verma et al., 2019; Yun et al., 2019; Harris et al., 2021). In addition, imbalance-aware mixup methods have been proposed to improve model robustness on imbalanced datasets (Chou et al., 2020; Galdran et al., 2021). Furthermore, selective mixup is introduced by selecting sample pairs for mixup based on specific criteria instead of randomly, and it improves the model’s ability for domain adaptation (Xu et al., 2020) and out-of-distribution (OOD) generalization (Yao et al., 2022).
@@
A recent line of research applies existing data augmentation methods in class-incremental learning to further improve accuracy. Most existing works simply combine experience replay methods with vanilla Mixup (Mi et al., 2020; Kumari et al., 2022), RandAugment (Zhang et al., 2022b), or a combination of existing data augmentation methods (Bang et al., 2021). Extending this line of work, (Qiang et al., 2023) improves the sampling and mixing factor strategies in Mixup to address distribution discrepancies between the imbalanced training set and the balanced test set in continual learning, which is similar to existing imbalance-aware mixup methods (Chou et al., 2020; Galdran et al., 2021). Another work (Kim et al., 2022) addresses class-incremental learning by using out-of-distribution detection methods that utilize image rotation as a data augmentation technique. Another line of research is using generative models to generate data for previous tasks (Shin et al., 2017; Gao and Liu, 2023; Kim et al., 2024a), but this approach incurs additional costs for continuously training generative models across tasks. In addition, there is a work (Zhu et al., 2021) to achieve comparable performance with experience replay methods without storing buffer data by utilizing class augmentation and semantic augmentation techniques. However, all these works simply use existing data augmentation techniques for further improvements and do not propose any novel data augmentation methods specifically designed for continual learning scenarios. In comparison, GradMix is a robust data augmentation method designed to minimize catastrophic forgetting in class-incremental learning. We leave more detailed comparisons with these works in Sec. C.
3. Preliminaries
@@
We set notations for class-incremental learning. We denote each task as $T_{l}=\{d_{i}=(x_{i},y_{i})\}_{i=1}^{k}$, where $x$, $y$, and $k$ represent the feature, the true label, and the number of samples, respectively. We assume disjoint sets of classes as tasks, where there are no overlapping classes between different tasks. The sequence of tasks is given by $T=\{T_{1},\ldots,T_{L}\}$, where $L$ is the total number of tasks. At the $l^{th}$ step, we consider $T_{l}$ as the current task and $\{T_{1},\ldots,T_{l-1}\}$ as previous tasks. Let the set of previous classes be $\mathbb{Y}_{p}$ and the set of current classes be $\mathbb{Y}_{c}$. We store a small subset of data for the previous tasks in a buffer to employ experience replay during training. We denote the subset of each previous task as $\mathcal{M}_{t}=\{d_{j}=(x_{j},y_{j})\}_{j=1}^{m}$, where $m$ is the number of samples. The buffer data for the previous tasks is then represented as $\mathcal{M}=\{\mathcal{M}_{1},\ldots,\mathcal{M}_{l-1}\}$. Since we cannot store large buffer data due to the possible problems of memory, efficiency, and privacy during continual learning, we assume the buffer size is small (i.e., $m\ll k$) (Chaudhry et al., 2019b).
@@
Catastrophic forgetting is the main challenge of continual learning, where the performance on previously-learned tasks degrades while adapting to the current task. The condition for the occurrence of catastrophic forgetting is theoretically derived using the gradients of previous and current data with respect to the model (Lopez-Paz and Ranzato, 2017; Chaudhry et al., 2019a; Aljundi et al., 2019b). At the $l^{th}$ step, let $G_{y}$ be a group of samples whose label is $y\in\mathbb{Y}_{p}$, and $T_{l}$ be the current task data used for training. We can approximate the average loss of $G_{y}$ after model training with $T_{l}$ by employing first-order Taylor series approximation as follows:
(1) $\tilde{\ell}(f_{\theta},G_{y})=\ell(f_{\theta}^{l-1},G_{y})-\eta\nabla_{\theta}\ell(f_{\theta}^{l-1},G_{y})^{\top}\nabla_{\theta}\ell(f_{\theta}^{l-1},T_{l}),$
[…]
@@
Mixup (Zhang et al., 2018) is a data augmentation method that generates new training samples by combining random pairs of existing samples of features and their labels. Given two arbitrary samples $d_{i}=(x_{i},y_{i})$ and $d_{j}=(x_{j},y_{j})$ from the dataset, a new mixed sample $\tilde{d}_{ij}=(\tilde{x}_{ij},\tilde{y}_{ij})$ is generated from the two original samples as follows:
(2) $\tilde{x}_{ij}=\lambda x_{i}+(1-\lambda)x_{j},\ \tilde{y}_{ij}=\lambda y_{i}+(1-\lambda)y_{j},$
@@
4. Catastrophic Forgetting in Mixup
In this section, we demonstrate that simply applying Mixup to experience replay methods in class-incremental learning can lead to greater catastrophic forgetting of previously learned classes. In this paper, we use a simple yet representative experience replay method, ER (Chaudhry et al., 2019b), as a basis method for class-incremental learning, although other methods could also be used.
4.1. Theoretical Analysis